OpenStudy (anonymous):

If a triangle has an angle of 45° and an angle of 100°, what is the third angle? A. 55° B. 35° C. 100° D. 45°

OpenStudy (calculusxy):

A triangle equals to 180 degrees. We can write an equation and the solve for the missing variables. 45+100+x=180

OpenStudy (anonymous):

180 is not a option

OpenStudy (calculusxy):

No I meant tht the total angle for a triangle is 180 degrees. That should help you find the misiing angle in the triangle.

OpenStudy (anonymous):

i agree 180 degrees but that what i am trying to find is the missing third angle finding it hard to come up with Thank for your help i appreciate it

OpenStudy (calculusxy):

An easy way would be to solve for the equation that I provided above 45+100+x=180 x is the misiing angle that we want to find. So add 45+100=145 then subtract 180-135=35 x=35. Answer is 35 degrees.

OpenStudy (calculusxy):

Sorry I meant 180-145=35 :)

OpenStudy (anonymous):

thank you i just couldn't figure it out ...

OpenStudy (calculusxy):

If you like my answer, you can give me a medal or be my fan :)